Hi,
I'm not able to find the area of a trapezoid with height 2ft and bases 1ft and 3ft.


Thanks

Found 2 solutions by checkley75, stanbon:
Answer by checkley75(3666)

Area of a trapazoid is height*(s1+s2)/2 where s1=side 1 7 s2=side2
2*(1+3)/2
2*4/2
2*2=4 ft^2 for the area

Answer by stanbon(57246)

find the area of a trapezoid with height 2ft and bases 1ft and 3ft.
----------
A = (1/2)h(b1+b2)
Area = (1/2)(2)(1+3)
Area = 1(4)
Area = 4 sq ft.
Cheers,
Stan H.